# The Analytic Stockwell Transform and its Zeros

### v2

@@ 3.2 Link with the zeros of the Hyperbolic Gaussian Analytic Function
Let $\beta>0$, $\psi_{\beta}\in L^{2}(\mathbb{R})$ be the Cauchy wavelet of parameter $\beta$ and $\varphi_{\beta}(t)=\psi_{\beta}(t)\mathrm{e}^{-2\pi\mathrm{i}t}$. Then, up to a conformal transform, the zeros of the Analytic Stockwell Transform with analysis window $\varphi_{\beta}$ of the white noise coincide in law with the zero set of the hyperbolic Gaussian Analytic Function of parameter $12\beta+$ 1. Consequently, the zero set of the Analytic Stockwell Transform of white noise is invariant under isometries of $\mathbb{D}$.
@@
Establishing a connection between time–frequency transforms and Gaussian Analytic Functions with isometry-invariant zero sets requires to consider the complex-valued white noise [5, Section 4], [6, 3, 24, 30]. Thus, all along this work, “white noise" will refer to a complex random variable. [^3] The interested reader can refer to [5, Section 3] for a study of the far more intricate case of real white noise, which will not be discussed in the present work.
@@
This proof follows closely the reasoning developed in [6, Theorem 4.3], linking the Cauchy wavelet transform of white noise to the Hyperbolic Gaussian Analytic Function. It is sketched here for the sake of completeness.
The first step consists in defining rigorously the continuous white noise adapted to the considered transform. To that aim, two main paths have bee, proposed, respectively described in [24, 5] and in [6], which is used in the following. Considering $\mathcal{H}=L^{2}(\mathbb{R})$ and the Hilbertian basis $\{f_{n}\}_{n\in\mathbb{N}}$ introduced in Lemma 1, applying [6, Proposition 3.3] yields a completion $\Theta$ of $L^{2}(\mathbb{R})$ in which the limit of the series
@@
with $\left(\zeta_{n}\right)_{n\in\mathbb{N}}$ a sequence of independent identically distribution standard Gaussian random variables, defines a Gaussian random variable with the characteristic function of the white noise. Note that, as emphasize in [6, Section 3] $\Theta$ depends on the basis, hence the importance of Lemma 1 deriving a basis suited to the Analytic Stockwell Transform.
The second step is to show that the Analytic Stockwell Transforms extends to $\Theta$, which then ensures that the transform of the continuous white noise is well-defined. To that aim, for $\beta>0$ fixed, let $\{\Psi_{n}\}_{n\mathbb{N}}$ be the sequence of analytic functions on $\mathbb{D}$ defined as
@@
Hence, the ($C_{\mathfrak{K}}$ revisited) condition of [6, Theorem 3.4] is satisfied and
@@
$\begin{split}\forall(x,\xi)\in\mathbb{R}\times\mathbb{R}_{+},\quad S_{\beta}f_{n}(x,\xi)&=\eta_{\beta}(x,\xi)\Psi_{n}(\vartheta(z))\\ &=\eta_{\beta}(x,\xi)\mathcal{T}f_{n}(\vartheta(z)),\quad z=x+\mathrm{i}\xi^{-1}\end{split}$ (35)
which extends to any $f\in L^{2}(\mathbb{R})$ by continuity of $\mathcal{T}$. Applying [6, Theorem 3.4] proves that $\mathcal{T}$ extends continuously to $\Theta$. Consequently, $\mathcal{T}\zeta$ is well-defined and it follows from the fact that $\eta_{\beta}$ is nonvanishing that the relation
$\forall(x,\xi)\in\mathbb{R}\times\mathbb{R}_{+},\quad S_{\beta}\zeta(x,\xi)=\eta_{\beta}(x,\xi)\mathcal{T}\zeta(\vartheta(z)),\quad z=x+\mathrm{i}\xi^{-1}$ (36)
@@
$S_{\beta}\zeta(x,\xi)=\eta_{\beta}(x,\xi)\mathsf{GAF}^{(2\beta+1)}(\vartheta(z)).$ (37)
@@ 4 Numerical Experiments and Spatial Statistics of Zeros
The present section goal is twofold. First, illustrations of the zero set of the Analytic Stockwell Transform are provided, enabling the reader to observe qualitatively the hyperbolic uniform spread of zeros in the Poincaré disk. Second, the characterization of zeros, established in Theorem 3, is supported by quantitative numerical evidence obtained by thorough comparison of the theoretical and empirical spatial statistics of the point process of zeros. A documented Python toolbox implementing the Analytic Stockwell Transform and performing the spatial statistics analysis of the zeros of the transform, thus enabling to reproduce all the experiments, plots and figures presented in Section 4 has been made publicly available by the authors. [^4] https://github.com/courbot/ast
@@ 4.1 Discrete Time–Frequency Analysis
corresponding to sampling $f$ in $[x_{\min},x_{\max}]$ at equally spaced time $x_{1},\ldots,x_{N}$ with resolution $\Delta_{x}=(x_{\max}-x_{\min})/(N-1)$. This yields a discrete signal $\mathbf{y}=(y_{1},\dots,y_{N})\in\mathbb{C}^{N}$. The continuous signal $f$ is then said to be sampled at frequency $1/\Delta_{x}$. Furthermore, in practice, the frequency variable needs to be discretized as well. In the present work, motivated by the connection with the Cauchy wavelet transform, the discrete Analytic Stockwell transform is computed on $M$ logarithmically spaced frequency channels: $\forall m\in\{1,\ldots,M\}$,
@@
where $\Delta_{\log_{2}\xi}$ is the log-frequency resolution, and the range $[\xi_{\min},\xi_{\max}]$ is chosen depending on the time–frequency pattern of interest, with the constraint of satisfying Nyquist-Shannon reconstruction theorem [38, Theorem 5.15].
@@
where $\zeta_{1},\ldots,\zeta_{N}$ are independent identically distributed Gaussian variables.
Proposition 1.
The discrete transform $\mathbf{S}_{\varphi_{\beta}}\boldsymbol{\zeta}$ is a discretization of the Analytic Stockwell Transform of continuous white noise $S_{\varphi_{\beta}}\zeta$ of Equation (36).
Proof.
@@
The demonstration relies on the orthogonality of the family $\{f_{n}\}_{n\in\mathbb{N}}$, defined in Lemma 1, used to construct the continuous white noise $\zeta$, and stems from the reasoning developed in [6, Section 5]. ∎
@@
Numerically, the discrete Analytic Stockwell Transform is computed in the Fourier domain, using a Riemann approximation of Equation (19) combined with Equation (9). Following [15, 3, 24, 5, 6, 30, 12] the localization of zeros is performed using the Minimal Grid Neighbors algorithm, introduced initially in the code [^5] https://perso.ens-lyon.fr/patrick.flandrin/zeros.html accompanying the seminal paper [15]. The method is described in details in [31, Section 4.3] in the context of standard time–frequency analysis, and can be applied as is to the discrete Analytic Stockwell Transform of Definition 5. A major advantage of the Minimal Grid Neighbors algorithm is that it benefits from a strong robustness, explained by its solid theoretical ground established recently [12].
@@
Figure 1 displays the modulus of the discrete Analytic Stockwell Transform of parameter $\beta$ such that $\alpha=2\beta+1=300$ computed over a realization of discrete white noise of length $N=4000$ with sampling frequency $\nu_{s}=400$ Hz, corresponding to a time resolution $\Delta_{x}=1/\nu_{s}$, and for $M=600$ frequency channels ranging from $\xi_{\min}=2^{-6}$ to $\xi_{\max}=2^{3.3}$, together with the associated zeros. The representation on the Poincaré disk, on the right of Figure 1, illustrates that, in the hyperbolic geometry in which the metric explodes as one gets closer to the border of the disk $\mathbb{D}$, the zeros spread very evenly. Similarly, the representation in the time–scale plane, on the left of Figure 1, shows the uniform spread in the hyperbolic upper-half plane.
@@
Figure 1: Log-modulus of the Analytic Stockwell Transform of white noise (background colormap) and its zero set (red dots) $\mathfrak{Z}^{(\alpha)}$. The discrete white noise is considered to span a time period of $x_{\max}-x_{\min}=10$ s and corresponds to a sampling frequency of $400$ Hz over. The parameter of the Cauchy wavelet $\beta$ is chosen such that $\alpha=2\beta+1=300$. Representation in the $(x,\xi^{-1})$-plane where the inverse frequency $\xi^{-1}$ is called the scale in reference to the related Cauchy wavelet transform (left). Representation in the Poincaré disk $\mathbb{D}$ parameterized by $\vartheta(x+\mathrm{i}\xi^{-1})$ for $(x,\xi^{-1})$ running over the same grid, where $\vartheta$ is the Cayley transform of Equation (10). The dotted circle represents the border of $\mathbb{D}$ (right).
@@ 4.2 Spatial Statistics
Point processes can be characterized by their $k$-points correlation functions [10, Section 5.4], describing the probabilistic interactions between the points of a realization of the point process, e.g., the short-range repulsion observed in the zero pattern of analytic time–frequency transforms [15, 24, 5, 6, 30].
@@
for any bounded compactly supported measurable map $\Psi:\Lambda^{k}\rightarrow\mathbb{C}$, and where $\mathrm{d}w$ denotes the metric in the ambient space, e.g., the hyperbolic metric $\mathsf{m}_{\mathbb{D}}$ in the Poincaré disk. In particular, $\rho_{1}$ is often called the intensity of the point process as, by definition, integrating it over a domain yields the expected number of points falling into. As for the two-point correlation function $\rho_{2}$, it is commonly normalized to get the pair-correlation function [28, Chapter 4], defined as
$\widetilde{g}(w_{1},w_{2})=\frac{\rho_{1}(w_{1},w_{2})}{\rho_{1}(w_{1})\rho_{1}(w_{2})},\quad w_{1},w_{2}\in\Lambda.$ (45)
@@
Remarkably, the joint intensities of zero sets of Gaussian Analytic Functions have closed-form expressions. Applying the Edelman-Kostlan formula [21, Section 2.4.1] [31, Proposition 16] yields the explicit expression [^6] The difference from the formula derived in [3, 24] comes from a difference in the normalization of the hyperbolic metric of Equation (12) and of the metric used in [3, 24]. of the intensity of the zeros of $\mathsf{GAF}_{\mathbb{D}}^{(\alpha)}$:
@@
To access the higher-order joint intensities, one can leverage the formulas involving determinants and permanents provided in [21, Corrolary 3.4.1], to get the following explicit formula for the pair correlation functions [3]
@@
where $s=1-r^{2}$. This theoretical expression is to be compared to the pair correlation function of the zeros of the Analytic Stockwell Transform of parameter $\beta=(\alpha-1)/2$ of white noise.
To provide quantitative numerical evidence supporting Theorem 3, the theoretical pair correlation $g^{(\alpha)}$ is compared to the empirical pair correlation function of the Analytic Stockwell Transform of parameter $\beta=(\alpha-1)/2$ of white noise. The need for accurate and robust estimators of empirical pair correlation functions resulted in a rich variety of spatial statistics estimation strategies with solid theoretical guarantees [28, Section 4.3]. Notably, edge correction techniques have been developed to handle the fact that in practice only a bounded window is observed [28, Section 4.3.3].
@@
Although derived in Euclidean geometry, the aforementioned spatial statistics framework [28] applies as is in the Poincaré disk, replacing the Euclidean metric and distance by their hyperbolic counterparts.
In the present work, the pair correlation function is estimated using the nonparametric estimator of [28, Section 4.3.5], the kernel being chosen as a step function of support size $h>0$. The number of zeros in the observed window being large enough, border effects are managed using the reduced-sampled estimate [28, Section 4.3.3]. Denoting by $\mathfrak{Z}^{(\alpha)}$ the zero set of the Analytic Stockwell transform of parameter $\beta=(\alpha-1)/2$ of white noise, the pair correlation function estimator writes
where $\mathds{1}_{\mathcal{A}}$ denotes the indicative function of the probabilistic event $\mathcal{A}$. To correct the bias induced by the observation of a bounded window the averaging over $z$ is performed only on the zeros which are far enough from the border of the observed hyperbolic window. Denoting $\mathfrak{B}^{(\alpha)}\subset\mathfrak{Z}^{(\alpha)}$ the zeros which are too close to the window’s border, represented as pink dots in Figure LABEL:sfig:isocontours, the averaging is performed on inner points $z\in\mathfrak{Z}^{(\alpha)}\setminus\mathfrak{B}^{(\alpha)}$, represented as dark red points in Figure LABEL:sfig:outer, while the counting runs on every points $w\in\mathfrak{Z}^{(\alpha)}$. The practical implementation of $\widehat{g}^{(\alpha)}$ amounts to building concentric rings of radius $r$ around each point in the zeros set, as illustrated in Figure LABEL:sfig:isocontours, counting the number of points in each hyperbolic ring, averaging it over the observed zeros, and finally normalizing by the theoretical first-intensity function provided in Equation (47).
The theoretical pair correlation function of $\mathsf{GAF}^{(\alpha)}_{\mathbb{D}}$, whose explicit expression is provided in Equation (48), is displayed as a dashed black line in Figure 3 for three values of $\alpha$. One remarks, going from the left to the right of Figure 3, that as $\alpha$ increases the short-range repulsion between zeros visible at the vicinity of $r=0$ weakens while the range of attraction between zeros, correspond to the bump above one, gets narrower and shifts toward $r=0$.
@@
To compare the expected and empirical behaviors of the zeros of the Analytic Stockwell Transform of white noise, $R=100$ realizations of the discrete white noise in dimension $N=4000$ are generated, their Analytic Stockwell Transforms are computed at time resolution $1/\nu_{s}$ with $\nu_{s}=4000$ Hz and accross $600$ frequency channels, yielding $R$ realizations of the point process $\mathfrak{Z}^{(\alpha)}$ constituted of their zeros; finally the $R$ the empirical pair correlation functions $\widehat{g}^{(\alpha)}$ are estimated from Equation (49), and averaged leading to the averaged empirical pair correlation function $\langle\widehat{g}^{(\alpha)}\rangle$ displayed as the solid blue line with star ticks in Figure 3 accompanied with their $5\%$ and $95\%$ quantiles. Empirical estimates show a good match with the theoretical pair correlation function for all three values of $\alpha$, although with a slightly larger $95\%$ credibility region for small $\alpha$. All observations made on the theoretic $g^{(\alpha)}$ remains true for its empirical estimate, providing a strong numerical evidence of the connection between the hyperbolic Gaussian Analytic zeros and the zeros of the Analytic Stockwell Transform of white noise.

### v3

@@ 3.2 Link with the zeros of the hyperbolic Gaussian analytic function
Let $\beta>0$, $\psi_{\beta}\in L^{2}(\mathbb{R})$ be the Cauchy wavelet of parameter $\beta$ and $\varphi_{\beta}(t)=\psi_{\beta}(t)\mathrm{e}^{-2\pi\mathrm{i}t}$. Then, up to a conformal transform, the zeros of the Analytic Stockwell Transform with analysis window $\varphi_{\beta}$ of white noise coincide in law with the zero set of the hyperbolic Gaussian analytic function of parameter $12\beta+$ 1. Consequently, the zero set of the Analytic Stockwell Transform of white noise is invariant under isometries of $\mathbb{D}$.
@@
Establishing a connection between time–frequency transforms and Gaussian analytic functions with isometry-invariant zero sets requires to consider the complex-valued white noise [6, Section 4], [7, 3, 25, 31]. Thus, all along this work, “white noise” will refer to a complex random variable. [^3] The interested reader can refer to [6, Section 3] for a study of the far more intricate case of real white noise, which will not be discussed in the present work.
@@
This proof follows closely the reasoning developed in [7, Theorem 4.3], linking the Cauchy Wavelet Transform of white noise to the hyperbolic Gaussian analytic function. It is sketched here for the sake of completeness.
The first step consists in defining rigorously the continuous white noise adapted to the considered transform. To that aim, two main paths have been proposed, respectively described in [25, 6] and in [7], the later being used in the following. Considering $\mathcal{H}=L^{2}(\mathbb{R})$ and the orthonormal basis $\{f_{n}\}_{n\in\mathbb{N}}$ introduced in Lemma 1, define a novel norm $\lVert\cdot\rVert_{\Theta}$ such that
$\forall f\in\mathcal{H},\quad\lVert f\rVert_{\Theta}^{2}=\sum_{n\in\mathbb{N}}\frac{1}{1+n^{2}}\lvert\langle f,f_{n}\rangle\rvert^{2}.$ (31)
Then, the completion $\Theta$ of $L^{2}(\mathbb{R})$ for the norm $\lVert\cdot\rVert_{\Theta}$ is such that the series
@@
with $\left(\zeta_{n}\right)_{n\in\mathbb{N}}$ a sequence of independent and identically distributed standard Gaussian random variables has a well-defined limit. Furthermore, by orthonormality of the basis $\{f_{n}\}_{n\in\mathbb{N}}$ [7, Proposition 3.3] ensures that Equation (32) defines a Gaussian random variable with the characteristic function of white noise. Note that, as emphasized in [7, Section 3] $\Theta$ depends on the orthonormal basis $\{f_{n}\}$, hence the importance of Lemma 1 deriving a basis suited to the Analytic Stockwell Transform.
The second step is to show that the Analytic Stockwell Transforms extends to $\Theta$, which then ensures that the transform of the continuous white noise is well-defined. To that aim, for $\beta>0$ fixed, let $\{\Psi_{n}\}_{n\mathbb{N}}$ be the sequence of analytic functions on $\mathbb{D}$ defined by
@@
Hence, the ($C_{\mathfrak{K}}$ revisited) condition of [7, Theorem 3.4] is satisfied and
@@
$\begin{split}\forall(x,\xi)\in\mathbb{R}\times\mathbb{R}_{+},\quad S_{\varphi_{\beta}}f_{n}(x,\xi)&=\eta_{\beta}(x,\xi)\Psi_{n}(\vartheta(z))\\ &=\eta_{\beta}(x,\xi)\mathcal{T}f_{n}(\vartheta(z)),\quad z=x+\mathrm{i}\xi^{-1}\end{split}$ (36)
which extends to any $f\in L^{2}(\mathbb{R})$ by continuity of $\mathcal{T}$. Applying [7, Theorem 3.4] proves that $\mathcal{T}$ extends continuously to $\Theta$. Consequently, $\mathcal{T}\zeta$ is well-defined and it follows from the fact that $\eta_{\beta}$ is nonvanishing that the relation
$\forall(x,\xi)\in\mathbb{R}\times\mathbb{R}_{+},\quad S_{\varphi_{\beta}}\zeta(x,\xi)=\eta_{\beta}(x,\xi)\mathcal{T}\zeta(\vartheta(z)),\quad z=x+\mathrm{i}\xi^{-1}$ (37)
@@
$S_{\varphi_{\beta}}\zeta(x,\xi)=\eta_{\beta}(x,\xi)\mathsf{GAF}^{(2\beta+1)}(\vartheta(z)).$ (38)
@@ 4 Numerical Experiments and Spatial Statistics of Zeros
The purpose of the present section is twofold. First, illustrations of the zero set of the Analytic Stockwell Transform of white noise are provided, enabling the reader to observe qualitatively the hyperbolic uniform spread of zeros in the Poincaré disk. Second, the characterization of the distribution of zeros, established in Theorem 3, is supported by quantitative numerical evidence obtained by thorough comparisons of the theoretical and empirical spatial statistics of the point process of zeros. A documented Python toolbox implementing the Analytic Stockwell Transform and performing the spatial statistics analysis of the zeros of the transform, thus enabling to reproduce all the experiments, plots and figures presented in Section 4 has been made publicly available by the authors. [^4] https://github.com/courbot/ast
@@ 4.1 Discrete Time–Frequency Analysis
corresponding to sampling $f$ in $[x_{\min},x_{\max}]$ at equally spaced time $x_{1},\ldots,x_{N}$ with resolution $\Delta_{x}=(x_{\max}-x_{\min})/(N-1)$. This yields a discrete signal $\mathbf{y}=(y_{1},\dots,y_{N})\in\mathbb{C}^{N}$. The continuous signal $f$ is then said to be sampled at frequency $1/\Delta_{x}$. Furthermore, in practice, the frequency variable needs to be discretized as well. In the present work, motivated by the connection with the Wavelet Transform, the discrete Analytic Stockwell Transform is computed on $M$ logarithmically spaced frequency channels: $\forall m\in\{1,\ldots,M\}$,
@@
where $\Delta_{\log_{2}\xi}$ is the log-frequency resolution, and the range $[\xi_{\min},\xi_{\max}]$ depends on the time–frequency pattern of interest [40, Theorem 5.15].
@@
where $\zeta_{1},\ldots,\zeta_{N}$ are independent Gaussian variables.
Remark 5.
@@
Following the same reasoning as the one developed in [7, Section 5], the Analytic Stockwell Transform of continuous white noise $S_{\varphi_{\beta}}\zeta$ of Equation (37) can be discretized consistently applying the discrete Stockwell transform of Definition 5 to discrete white noise. This is notably due to orthogonality of the family $\{f_{n}\}_{n\in\mathbb{N}}$, defined in Lemma 1, used to construct the continuous white noise $\zeta$.
@@
Numerically, the discrete Analytic Stockwell Transform is computed in the Fourier domain, using a Riemann approximation of Equation (19) combined with Equation (9). Following [16] the localization of zeros is performed using the Minimal Grid Neighbors algorithm, introduced initially in the Euclidean setting in the code [^5] https://perso.ens-lyon.fr/patrick.flandrin/zeros.html accompanying the seminal paper [16], and which were then used in Euclidean [6, 3], hyperbolic [25] and spherical geometries [7, 31]. The method is described in details in [32, Section 4.3] in the context of standard time–frequency analysis, and is applied as is to the discrete Analytic Stockwell Transform of Definition 5.
Remark 6.
Recently an Adaptive Minimal Grid Neighbors algorithm has been derived in Euclidean geometry, benefiting from a solid theoretical ground and strong robustness [13]. Intensive numerical experiments has shown that both the Adaptive and standard Minimal Grid Neighbors algorithms perform equivalently and significantly outperform thresholding strategies [13, Section 2.3]. Extending this theoretical analysis to the hyperbolic geometry is beyond the scope of the present work, though it constitutes a very interesting line of research for future work.
@@
Figure 1 displays the modulus of the discrete Analytic Stockwell Transform of parameter $\beta$ such that $\alpha=2\beta+1=300$ computed over a realization of discrete white noise of $N=4000$ points with sampling frequency $\nu_{s}=400$ Hz, corresponding to a time resolution $\Delta_{x}=1/\nu_{s}$, and for $M=600$ frequency channels ranging from $\xi_{\min}=2^{-6}$ to $\xi_{\max}=2^{3.3}$, together with the associated zeros. The representation on the Poincaré disk, on the right of Figure 1, illustrates that, in the hyperbolic geometry in which the metric explodes as one gets closer to the border of the disk $\mathbb{D}$, the zeros spread very evenly. Similarly, the representation in the time–scale plane, on the left of Figure 1, shows the uniform spread in the hyperbolic upper-half plane.
@@
Figure 1: Log-modulus of the Analytic Stockwell Transform of white noise (background colormap) and its zero set (red dots) $\mathfrak{Z}^{(\alpha)}$. The discrete white noise is considered to span a time period of $x_{\max}-x_{\min}=10$ s with a sampling frequency of $400$ Hz. The parameter of the Cauchy wavelet $\beta$ is chosen such that $\alpha=2\beta+1=300$. (Left) Representation in the $(x,\xi^{-1})$-plane where the inverse frequency $\xi^{-1}$ is called the scale in reference to the related Cauchy Wavelet Transform. (Right) Representation in the Poincaré disk $\mathbb{D}$ parameterized by $\vartheta(x+\mathrm{i}\xi^{-1})$ for $(x,\xi^{-1})$ running over the same grid, where $\vartheta$ is the Cayley transform of Equation (10). The dotted circle represents the border of $\mathbb{D}$.
@@ 4.2 Spatial Statistics
Point processes in a general metric space can be characterized by their $k$-points correlation functions [11, Section 5.4], describing the probabilistic interactions between the points of a realization of the point process, e.g., the short-range repulsion observed in the zero pattern of analytic time–frequency transforms [16, 25, 6, 7, 31]. As emphasized in [29, Appendices B,C], spatial statistics tools can be derived from a general metric on any ambient space [11]. As this work focuses on the zeros of the Stockwell Transform lying in the Poincaré disk equipped with the hyperbolic distance introduced in Definition 3, all tools and mathematical results in the remaining of the paper are formulated directly in the hyperbolic geometry framework.
@@
for any bounded compactly supported measurable map $\Psi:\Lambda^{k}\rightarrow\mathbb{C}$, and where $\mathrm{d}w$ denotes the metric in the ambient space, e.g., the hyperbolic metric $\mathrm{d}\mathsf{m}_{\mathbb{D}}$ in the Poincaré disk. In particular, $\rho_{1}$ is often called the intensity of the point process as, by definition, integrating it over a domain yields the expected number of points falling into. As for the two-point correlation function $\rho_{2}$, it is commonly normalized to get the pair-correlation function [29, Chapter 4], defined as
$\widetilde{g}(w_{1},w_{2})=\frac{\rho_{2}(w_{1},w_{2})}{\rho_{1}(w_{1})\rho_{1}(w_{2})},\quad w_{1},w_{2}\in\Lambda.$ (46)
@@
Remarkably, the joint intensities of zero sets of Gaussian analytic functions have closed-form expressions. Applying the Edelman-Kostlan formula [22, Section 2.4.1] [32, Proposition 16] yields the explicit expression [^6] The difference from the formula derived in [3, 25] comes from a difference in the normalization of the hyperbolic metric of Equation (12) and of the metric used in [3, 25]. of the intensity of the zeros of $\mathsf{GAF}_{\mathbb{D}}^{(\alpha)}$:
@@
To access the higher-order joint intensities, one can leverage the formulas involving determinants and permanents provided in [22, Corrolary 3.4.1], to get the following explicit formula for the pair correlation functions [3]
@@
where $s=1-r^{2}$. By Theorem (3), Equation (49) corresponds both to the theoretical pair correlation function of the zeros of $\mathsf{GAF}^{(\alpha)}_{\mathbb{D}}$ and of the zeros of the Analytic Stockwell Transform of parameter $\beta=(\alpha-1)/2$ of white noise. For reference, $g^{(\alpha)}$ is displayed as a dashed black line in Figure 3 for three values of $\alpha$. One remarks, going from the left to the right of Figure 3, that as $\alpha$ increases the short-range repulsion between zeros visible at the vicinity of $r=0$ weakens while the range of attraction between zeros, correspond to the bump above one, gets narrower and shifts toward $r=0$. To provide quantitative numerical evidence showing practical relevance of Theorem 3, the theoretical expressions of both the first intensity and the pair correlation function are to be compared to the empirical spatial statistics of the zeros of the Analytic Stockwell Transform of parameter $\beta=(\alpha-1)/2$ of white noise.
The need for accurate and robust estimators of empirical first intensity and pair correlation functions resulted in a rich variety of spatial statistics estimation strategies with solid theoretical guarantees [29, Section 4.3]. Notably, edge correction techniques have been developed to handle the fact that in practice only a bounded window is observed [29, Section 4.3.3].
First of all, to assess the ability of the hyperbolic counterpart of Minimal Grid Neighbor algorithm to correctly detect zeros in hyperbolic geometry with the discretization strategy described in Definition 5, the first intensity $\rho_{1}^{(\alpha)}$ of the zeros of the discrete Stockwell Transform of white noise of parameter $\beta=(\alpha-1)/2$ is estimated by counting the number of zeros in a growing hyperbolic disk centered at zero and of increasing hyperbolic radius $r^{\prime}$ and dividing it by the hyperbolic area $\mathcal{A}(r^{\prime})$ of this disk, obtained integrating the hyperbolic metric (12),
$\hat{\rho}_{1}^{(\alpha)}=\frac{1}{\mathcal{A}(r^{\prime})}\sum_{z\in\mathfrak{Z}^{(\alpha)}}\mathds{1}_{\mathsf{d}_{\mathbb{D}}(z,0)<r^{\prime}},\quad\mathcal{A}(r^{\prime})=4\pi\sinh(r^{\prime}/2)^{2}$ (50)
where $\mathfrak{Z}^{(\alpha)}$ denotes the zero set of the Analytic Stockwell transform of parameter $\beta=(\alpha-1)/2$ of white noise [^7] Since hyperbolic and pseudo-hyperbolic distances are in one-to-one correspondence the above computations could be implemented equivalently replacing $\mathsf{d}_{\mathbb{D}}$ by $\mathsf{p}_{\mathbb{D}}$ and $r^{\prime}$ by $r=\tanh(r^{\prime}/2)$ according to (13). and $\mathds{1}_{\mathcal{A}}$ denotes the indicative function of the probabilistic event $\mathcal{A}$.
Then, the pair correlation function is estimated using the nonparametric estimator of [29, Section 4.3.5], the kernel being chosen as a step function of support size $h>0$. The number of zeros in the observed window being large enough, border effects are managed using the reduced-sampled estimate [29, Section 4.3.3]. The pair correlation function estimator writes
where $\mathfrak{Z}^{(\alpha)}$ is the zero set of the Analytic Stockwell transform of parameter $\beta=(\alpha-1)/2$ of white noise. To correct the bias induced by the observation of a bounded window the averaging over $z$ is performed only on the zeros which are far enough from the border of the observed hyperbolic window. Denoting $\mathfrak{B}^{(\alpha)}\subset\mathfrak{Z}^{(\alpha)}$ the zeros which are too close to the window’s border, represented as pink dots in Figure LABEL:sfig:isocontours, the averaging is performed on inner points $z\in\mathfrak{Z}^{(\alpha)}\setminus\mathfrak{B}^{(\alpha)}$, represented as dark red points in Figure LABEL:sfig:outer, while the counting runs on every points $w\in\mathfrak{Z}^{(\alpha)}$. The practical implementation of $\widehat{g}^{(\alpha)}$ amounts to building concentric rings of pseudo-hyperbolic radius $r$ around each point in the zeros set, as illustrated in Figure LABEL:sfig:isocontours, counting the number of points in each pseudo-hyperbolic ring, which by (13) are also hyperbolic rinds, averaging it over the observed zeros, and finally normalizing by the theoretical first-intensity function provided in Equation (48).
@@
It is worth insisting on the fact that, although mostly used in Euclidean spaces, the aforementioned spatial statistics framework applies in any general metric space [29, Appendices B,C]. The customization to hyperbolic geometry then reduces to replacing the Euclidean metric and distance by their hyperbolic counterparts. This affects the shapes of the disks and the identification of outer points, as can be seen on Figure 2, but not the theoretical results, proven in a general setting [29, Appendices B,C], [11, Chapters 6,7]. Notably, the edge correction implemented in Equation (51) ensures the unbiasedness of the pair correlation function estimator regardless of the geometry.
@@
To compare the expected and empirical behaviors of the zeros of the Analytic Stockwell Transform of white noise, $R=100$ realizations of the discrete white noise with $N=4000$ time stamps are generated. Their discrete Stockwell Transforms are computed at time resolution $1/\nu_{s}$ with $\nu_{s}=4000$ Hz and across $600$ frequency channels. The Minimal Grid Neighbors algorithm then yields $R$ realizations of the point process $\mathfrak{Z}^{(\alpha)}$ each consisting in the zeros of a discrete Stockwell Transform of white noise. Finally the $R$ empirical first intensities $\widehat{\rho}_{1}^{(\alpha)}$ and pair correlation functions $\widehat{g}^{(\alpha)}$ are estimated from Equation (51), and averaged leading to the averaged empirical first intensity $\langle\widehat{\rho}_{1}^{(\alpha)}\rangle$ and pair correlation function $\langle\widehat{g}^{(\alpha)}\rangle$ displayed respectively in Figure 3 left and right columns as solid blue lines with star ticks, accompanied with their $5\%$ and $95\%$ quantiles.
Provided that the hyperbolic disk used to estimate the first intensity is large enough, the empirical estimates $\widehat{\rho}_{1}^{(\alpha)}$ are very close to the expected value provided in Equation (48). This means that the Minimal Grid Neighbors algorithms, applied to the discrete Stockwell Transform of Definition 5, detects accurately enough the zeros of the transform so that the first order statistics are in very good agreement with the known statistics of the zeros of the hyperbolic Gaussian analytic function. This, despite the fact that the discretization of the Stockwell transform, regular in time and in frequency, is suboptimal from an hyperbolic geometry point of view. Furthermore, empirical estimates $\widehat{g}^{(\alpha)}$ show a good match with the theoretical pair correlation function for all four values of $\alpha$, although with a slightly larger $95\%$ credibility region for small $\alpha$. Overall, the first and second order spatial statistics of the zeros of the discrete Stockwell Transform of white noise resemble very much the theoretical statistics of the zeros of the hyperbolic Gaussian analytic function.
Remark 8.
Note that for all the numerical experiments presented in this paper, the sampling frequency is kept the same. Furthermore, the proposed implementation of the discretized Analytic Stockwell Transform [^8] https://github.com/courbot/ast makes use of the Fast Fourier Transform algorithm and thus the frequencies on which the discrete Stockwell Transform is evaluated is fixed as well. As a consequence, both the time and frequency bin sizes are independent of $\alpha$ which controls the size of the support of the analysis window $\varphi_{\beta}$ through $\beta=(\alpha-1)/2$. Consequently for very low values of $\alpha$, the time and frequency bin sizes of the discrete Analytic Stockwell Transform turn out to be inappropriate to capture correctly the time–frequency information encoded in the continuous transform. This affects the numerical retrieval of zeros and makes it less accurate which explains why the difference between the theoretical and numerical estimates of the pair correlation function increases as $\alpha$ decreases as illustrated in Figure 3.
These experiments show that the empirical zeros of the discrete Stockwell Transform of white noise behaves statistically very similarly compared to the zeros of the hyperbolic Gaussian analytic function. This opens the way for leveraging theses statistics to perform practical signal processing tasks: such as detection or denoising following the recent works [16, 3, 25, 6, 31].